an=3an−1+n2−3,n≥,a0=1


1
Expert's answer
2022-01-24T14:34:12-0500
"a_n=3a_{n-1}+n^2-3, n\geq0, a_0=1"


The homogeneous part of this recurrence relation is;


"a_n=3a_{n-1}"

The characteristics equation for this is;


"r=3"


Thus, the solution for the homogeneous part is;


"a_n=\alpha3^n"


The complementary solution is:


"a_n^{(c)}= \alpha 3^n"

The non homogeneous part of the relation is "n^2-3"

In this case, the particular solution is of the form:


"a_n^{(p)}=p_2n^2+p_1n+p_0"

Substituting into the original relation, we have that:


"p_2n^2+p_1n+p_0=3(p_2(n-1)^2+p_1(n-1)+p_0)+n^2-3\\
p_2(-2n^2+6n-3)+p_1(-2n+3)-2p_0=n^2-3\\
-2p_2n^2+(6p_2-2p_1)n+(-3p_2+3p_1-2p_0)=n^2-3\\"

We now have that;

"-2p_2=1 \implies p_2=-\frac{1}{2}\\
6p_2-2p_1=0 \implies p_1=3p_2 \implies p_1=-\frac{3}{2}\\
-3p_2+3p_1-2p_0=-3 \implies -2p_0=0 \implies p_0=0"

Thus the particular solution is


"a_n^{(p)}=-\frac{1}{2}n^2-\frac{3}{2}n"

Hence, the general solution which is "a_n^{(c)}+a_n^{(p)}" is:


"a_n=\alpha 3^n-\frac{1}{2}n^2-\frac{3}{2}n"

Substitute "a_0=1" to get the value of "\alpha" . We have that;


"a_0=\alpha 3^0 \implies \alpha=1"

Hence the general solution of the recurrence relation is:


"a_n=3^n-\frac{1}{2}n^2-\frac{3}{2}n, n\geq 0"